OpenStudy (anonymous):

the fastest a human has ever run is 27 miles per hour. How many miles per minute did the human run.

geerky42 (geerky42):

\[27\dfrac{miles}{hour}\] We know that \(60~min=hour\) So \(27\dfrac{miles}{hour} = \dfrac{27}{60}\dfrac{miles}{min} = \cdots ~?~\dfrac{miles}{min}\)